A student applies to two jobs, one at Google and one at Apple. The probability of being hired at Google is 0.2, and independently, the probability of being hired at Apple is 0.6.
(a)What is the probability of receiving job offers from both companies?
(b) What is the probability of receiving job offers from at least one of the companies?
(c) What is the probability of receiving exactly one job offer? (d) What is the probability of not receiving any job offers?

Answer by ikleyn(52780)

.

(a)  P = P(G)*P(A) = 0.2*0.6 = 0.12.


(b)  P = P(G) + P(A) - P(G)*P(A) = 0.2 + 0.6 - 0.2*0.6 = 0.68.    


(c)  P = (P(G) - P(G)*P(A)) + (P(A) - P(G)*P(A)) = (0.2-0.12) + (0.6-0.12) = 0.56.         


(d)  P = the complement to the probability of the n.(b) = 1 - 0.68 = 0.32.